The measure of one angle of a triangle is twice the measure of the smallest angle. The measure of the third angle is 44 degrees more than 5 times the measure of the smallest angle. Find the measure of the largest angle.
Answer by ccs2011(207)

Let x be the smallest angle.
The sum of interior angles of a triangle equal 180.
x + 2x + (5x + 44) = 180
8x + 44 = 180
8x = 136
x = 17
Largest angle = 5x +44
--> 5(17) + 44 = 85 + 44 = 129
Measure of largest angle is 129 degrees.